"Q4) An analyst gathered the following information for a stock and market parameters: stock beta = 1.38; expected return on the Market = 10.60%; expected return on T-bills = 2.80%; current stock Price = $5.43; expected stock price in one year = $11.88; expected dividend payment next year = $2.30. Calculate the Expected return for this stock
Answer #1

Required return for the stock=risk free rate+beta*(market return-risk free rate)=2.80%+1.38*(10.60%-2.80%)=13.56400%

Expected return for the stock is given as=(D1+P1)/P0-1=(11.88+2.30)/5.43-1=161.141%